# Implementing Fairness in AI Classification: The Role of Explainability

### v2

@@ 5.4 Sample Weights vs Optimization under Constraint
For example, the new weight $w_{2}(S_{1})=0.8$ of any 17-29-year old individual $i$ in the initial observation 3 means that if a 17-29-year old $i$ has a true label of ‘earning less than \$50,000’ ($y_{i}=0$), predicting her as earning over \$50,000 ($\widehat{y_{i}}=1$) amounts to 8 times the cost of any prediction error for 29-37 years old. The FairDream-style classifier is, therefore, strongly incentivized to give predictions in accordance with the true label $y_{i}\in\{0,1\}$. As a consequence, the model becomes more accurate across age groups, by predicting $y_{i}=1$ given their true labels. Which makes the model tend towards Equalized Odds (equality of true and false positive rates).
@@ 6 Choosing a Fairness Metric conditioned on True Labels
We provided a computational and mathematical explanation for the deviation between the initial fairness objective (Demographic Parity) and the one actually achieved by FairDream (Equalized Odds). In this section we propose a more normative discussion of the choice between these two fairness objectives. As announced in the introduction, Equalized Odds is also commonly opposed to various notions of predictive parity, understood either in terms of equal calibration ([Pleiss et al., 2017, Barocas et al., 2023]), or in terms of equal precision relative to a threshold ([Dieterich et al., 2016, Mayson, 2018]). So first, we start with some data relative to FairDream’s performance on calibration, and we explain why, on a normative basis, calibration for us does not constitute a relevant criterion. The real issue, according to us, is whether the achievement of Equalized Odds remains too conservative compared to the achievement of Demographic Parity.
6.1 Calibration and Precision
@@
One common view of fairness is found in the idea of predictive parity between groups. Predictive parity can mean that the same fraction of true positives is predicted between groups at identical risk scores, or it can mean that the same fraction of true positives is predicted between groups relative to a classification threshold. The first notion corresponds to the statistical notion of calibration of a classifier, whereas the second corresponds to the notion of statistical precision of a classifier. Both notions are sometimes indistinctly referred to under the term “calibration” in the literature (see [Mayson, 2018, Long, 2021]), although strictly speaking they differ. Both notions have also been invoked in defense of COMPAS against ProPublica’s charges of racial bias. In their response to ProPublica, Northpointe argued that at the cutoff chosen by ProPublica to show Unequal Odds between Whites and Blacks, the statistical precision of the COMPAS algorithm was roughly the same between both groups. Moreover, it was shown that at each risk score, the proportion of rearrest between groups was roughly the same (see [Corbett-Davies et al., 2017, Figure 2]),141414[Dieterich et al., 2016] produce very closely related data (see their tables A1 and A2), but they look at predictions above all possible risk scores, which is not strictly speaking calibration. this time evidencing calibration between groups.
@@
The upshot is that Equal Precision is neither a necessary, nor a sufficient condition on fairness. In the case we just presented, one might respond that Precision in the “Low” judgments is not constant across the groups, even as precision on the “High” judgments is. But the problem is that to ask for Equal Precision on both labels is tantamount to asking for Equalized Odds, which would undercut Long’s argument. This is not to say that Precision does not matter to FairDream. As explained in the previous section (see fn. 12), we conducted our benchmark so as to maximize the trade-off between Precision and Recall. From Table 3, we also see that the gap in PR-AUC diminishes in FairDream across groups compared to the baseline. But since this is not a criterion we set a priori, no more than ROC-AUC, we do not build any conclusion on this fact.
@@
In summary, therefore, whether on a descriptive basis, or on a normative basis, Equal Calibration does not end up as a decisive criterion in our evaluation of FairDream’s correction, and neither does Equal Precision. In what follows, therefore, we return to the opposition between Equalized Odds and Demographic Parity as opposite requirements.
6.2 Fairness from Accuracy
In matters of fact, the case for conditional fairness metrics can be buttressed by general epistemological considerations. A user trying to narrow the gap between overall positive rates made on age groups by our FairDream classifier (16% for 17-29 years old and 59% for 29-37 years old, Cf. Table 2) may be viewed as falling prey to a “base rate neglect” ([Tversky et al., 1982]). It is a common mistake when interpreting statistics to overlook the base rates. In the present case, the base rate, or the ratio of actual positives to the whole population, is radically different in each age group. Less than one in six younger individuals actually earns over \$50,000, while four in six older actually earn over \$50,000.171717The base rates are 547/3616=15% for 17-29 years population, versus 2400/3666=65% for 29-37 years population (Cf. Table 2).
6.3 Always prefer Equalized Odds?
7 Conclusion
In this paper we implemented an algorithmic fairness tool to detect and correct for disparities in classification, the package FairDream. The practical goal of this package ‘is to help lay users implement their vision of fairness, in order to come up with socially optimal decisions. Our study of the specific properties of FairDream raised a puzzle and an explainability issue: how can an algorithm designed to enforce Demographic Parity end up converging on a different fairness goal, namely Equalized Odds?
First, we proposed an explanation for this difference by looking at the specifics of each algorithm. GridSearch  incorporates the fairness constraint as an additional cost inside the cost function of the model. Due to the fairness cost, this family of methods is more likely to deviate from the true labels (increasing false positives), in order to achieve Demographic Parity. On the other hand, FairDream introduces a differentiated attention to harmed individuals inside the cost function of the model. Therefore, it increases accuracy on harmed groups (through enhanced true positives). By basing fairness on accuracy, FairDream-type methods are thus bound to make the overall positive rates converge towards their base rates. As a result, they will rarely achieve Demographic Parity. But the method, we have also argued, also leads us to the conclusion that Equalized Odds is more relevant than either Equal calibration or Equal Precision in relation to fairness intervention.
Our study of FairDream also led us to a normative conclusion about how to approach fairness. By choosing a conditional metric, the user implicitly commits to the idea that there is sufficient agreement on the transparency of labels and on the shape of their distribution; by selecting an unconditional metric, one commits to the idea that the labels themselves may have introduced unwanted disadvantages, or that the input distribution must be corrected rather than replicated. In the case of FairDream, the default is therefore set to a correction method that may be viewed as overly prudent, but this is as it should be, since as we argued, more revisionary methods need to be carefully justified and cannot be applied lightly.
Finally, our investigation of FairDream allows us to draw a more general lesson concerning the explainability of AI models. Prima facie, FairDream may be seen as operationalizing a simple interventionist fairness metric, namely Demographic Parity. In practice, however, it produces results that are significantly different and that enforce Equalized Odds. As we have argued, the corrections produced by FairDream can be rationalized and normatively justified, but this involves an explanation of the reweighting method used by the algorithm, and a contrastive look at nearby systems governed by similar correction mechanisms that produce different outcomes. We conclude that AI classification systems should preferably be released with an indication of these features: which goal they seek to operationalize, how they actually perform relative to this goal, and finally how they perform on alternative data, and relative to minimally modified variants on the same data. Making this information available dispels arbitrariness in classification, and is a step toward more trust and more control on the side of the users.
Appendix
A FairDream’s Reweighting
We provide here more details on how the sample weights are generated during the in-processing correction of FairDream. The differentiated weights of error are established in a deterministic but simple way, to grant user’s understanding on the process of correction for fairness. For a FairDream model $n$in $[1,5]$ of five models to be trained in competition, the new weight of error for each group ${S_{k}}$ is computed that way:
B Benchmark on Census - models and features
Max Gap between groups OPR FPR TPR ROC-AUC PR-AUC Calibration All
GridSearch 7 9 10 11 12 8 57
FairDream 9 7 6 3 2 8 35
Each model received one point (of penalty) when it achieved the highest gap between the groups with the minimal and maximal scores. In Table 6, lower scores indicate better results on that count. To verify the convergence property of FairDream to Equalized Odds, the table also indicates the number of times a model was worse than the other at equalizing true positive rates, and false positive rates, across the relevant features.

### v3

@@ 5.4 Sample Weights vs Optimization under Constraint
For example, the new weight $w_{2}(S_{1})=0.8$ of any 17-29-year old individual $i$ in the initial observation 3 means that if a 17-29-year old $i$ has a true label of ‘earning less than \$50,000’ ($y_{i}=0$), predicting her as earning over \$50,000 ($\widehat{y_{i}}=1$) amounts to 8 times the cost of any prediction error for 29-37 years old (weight of 0.1). The FairDream-style classifier is, therefore, strongly incentivized to give predictions in accordance with the true label $y_{i}\in\{0,1\}$. As a consequence, the model becomes more accurate across age groups, by predicting $y_{i}=1$ given their true labels. Which makes the model tend towards Equalized Odds (equality of true and false positive rates).
@@ 6 Choosing a Fairness Metric conditioned on True Labels
We provided a computational and mathematical explanation for the deviation between the initial fairness objective (Demographic Parity) and the one actually achieved by FairDream (Equalized Odds). In this section and the next one, we propose a more normative discussion of the choice between these two fairness objectives. In this section, we first discuss whether Equalized Odds is too conservative from the perspective of a policy maker. In the next, we look at the constraint from an individual point of view, this time from the point of view of an individual in the situation of applying for a loan.
6.1 Fairness from Accuracy
In matters of fact, the case for conditional fairness metrics can be buttressed by general epistemological considerations. A user trying to narrow the gap between overall positive rates made on age groups by our FairDream classifier (16% for 17-29 years old and 59% for 29-37 years old, Cf. Table 2) may be viewed as falling prey to a “base rate neglect” ([Tversky et al., 1982]). It is a common mistake when interpreting statistics to overlook the base rates. In the present case, the base rate, or the ratio of actual positives to the whole population, is radically different in each age group. Less than one in six younger individuals actually earns over \$50,000, while four in six older actually earn over \$50,000.181818The base rates are 547/3616=15% for 17-29 years population, versus 2400/3666=65% for 29-37 years population (Cf. Table 2).
6.2 Always prefer Equalized Odds?
7 Conditional dependencies
Finally, in this section, we examine FairDream’s convergence towards Equalized Odds from the view point of an individual in the situation of being lent a loan. Has the FairDream model — whose predictions are shown in Table 2 — actually become fairer than the initial model shown in Table 1? We think the answer is positive, but it involves a careful look at dependencies between the relevant variables. The results produced by FairDream bear a resemblance to Simpson’s reversal paradox in that regard ([Pearl, 2009, Fitelson, 2017]), and we propose to use it as a guide toward our conclusion.
Let us first establish the analogy with Simpson’s paradox. On the one hand, the new model appears to strongly favor clients aged between 29 and 37 compared to the 16-29 group. On the other hand, when we split each group into those truly earning over 50k and others, this relation reverses within each group. That is, between the two age groups, the probability of being classified as earning more than 50k is equalized with FairDream. Formally, we have the following relations (where $A=1$ stands for membership in the 29-37 group; $B=1$ membership in the 16-29; $Y=1$ truly earning over 50k):
$p(\hat{Y}=1\mid A=1)\gg p(\hat{Y}=1\mid B=1)$
yet,
$p(\hat{Y}=1\mid A=1\ \&\ Y=1)\approx p(\hat{Y}=1\mid B=1\ \&\ Y=1)$
$p(\hat{Y}=1\mid A=1\ \&\ Y=0)\approx p(\hat{Y}=1\mid B=1\ \&\ Y=0)$
Technically, Simpson’s paradox requires a strict inversion within the partition groups (here, our groups of true labels). [^24] We are indebted to XXX for helping us clarify this matter. For it to qualify as a genuine Simpson’s paradox, it would have to be the case that, for every true label $i\in\{0,1\}$,
$p(\hat{Y}=1\mid A=1\ \&\ Y=i)\;\leq\;p(\hat{Y}=1\mid B=1\ \&\ Y=i).$
The direction of this inequality is not respected by the FairDream model, since we have:
$p(\hat{Y}=1\mid A=1\ \&\ Y=1)=88\%\;<\;p(\hat{Y}=1\mid B=1\ \&\ Y=1)=89\%$
$p(\hat{Y}=1\mid A=1\ \&\ Y=0)=4\%\;>\;p(\hat{Y}=1\mid B=1\ \&\ Y=0)=3\%.$
Among the younger clients who were formerly disadvantaged, FairDream now selects a slightly larger proportion of individuals than among their older counterparts within one group ($Y=1$), but not within the other ($Y=0$). Yet, this difference amounts to only 1%. Because of that, the fairness intervention occasioned by FairDream is a case of near-Simpson reversal, and we can draw on analyses of Simpson’s paradox to tease apart two readings of the fairness intervention.
Consider a 16-29 year-old who is wondering about their chances of being granted a loan, assuming it depends on an adequate recognition of their income. They learn that their chances (16%) are significantly lower than for a 29-37 year old (59%) even after correction by FairDream. Prima facie, this seems to support the following conditional statement:
[…]
. If I were in the 29-37 age group, my chances of getting a loan would be significantly higher than they are.
But this is a very weak case for lack of fairness. For this conditional does not necessarily keep constant the person’s assets under the counterfactual supposition changing the person’s age. Our argument is that either of the following conditionals gives a much better evaluation for fairness or lack of fairness, depending on the person’s income, which they are taken to know:
[…]
. If I were in the 29-37 age group, and I earned over 50k (as I do), my chances of getting a loan would be significantly higher than they are.
[…]
. If I were in the 29-37 age group, and I earned less than 50k (as I do), my chances of getting a loan would be significantly higher than they are.
Note that these two conditionals would come out true of the uncorrected model of Table 1, but they come out false of the corrected model proposed by FairDream. In other words, FairDream achieves near conditional independence of the predictor’s decision relative to the age variable, unlike the baseline model.
Using $a$ for “if I were in the 29-37 age group”, $b$ for “I earn over 50k” and $c$ for “my chances of getting a loan are higher than they are”, we can represent the previous sentences more succinctly as follows:
[…]
.˙ex:plain $a\to c$ True .̱ $a\wedge b\to c$ False .̧ $a\wedge\neg b\to c$ False
As it turns out, this triad of judgments is inconsistent in standard treatments of counterfactual reasoning. In Stalnaker’s or Lewis’s semantics for counterfactual conditionals, $a\to c$ entails the disjunction $((a\wedge b)\to c)\vee((a\wedge\neg b)\to c)$. The reason is that the closest world (or worlds) where I am in the 29-37 group is either a world where I earn over 50k, or a world where I earn under 50k. Which world it is is determined by which of these two options holds in the actual world. We consider this constraint to be sound, and to be reflected in the constraint of Equalized Odds. By contrast, the assertion of 7 corresponds to a counterfactual assumption in which one’s income can vary with one’s age. But this no longer corresponds to a ceteris paribus revision of one’s current situation. A more adequate paraphrase for 7 would therefore be: “if I were a person drawn randomly from the 29-37 age group, then on average I would stand higher chances of getting a loan that if I were a person drawn randomly from the 16-29 age group”. This is true, but this is not an adequate basis to conclude that the intervention produced by FairDream would be biased or unfair. For the base rates already show a similar imbalance regarding income across the groups, prior to any intervention.
In summary, we thus take the correction at play in FairDream to implement a notion of fairness that comes closer to ceteris paribus reasoning than the GridSearch methods we discussed. This is, ultimately, why we think of Equalized Odds as encapsulating a more adequate notion of fairness than Demographic Parity. [^25] We are indebted to the editors for prompting this discussion.
8 Conclusion
In this paper we implemented an algorithmic fairness tool to detect and correct for disparities in classification, the package FairDream. The practical goal of this package ‘is to help lay users implement their vision of fairness’, in order to come up with socially optimal decisions. Our study of the specific properties of FairDream raised a puzzle and an explainability issue: the algorithm is presented as aiming to equalize the percentage of individuals satisfying a certain condition across different groups – the direct interpretation of which is Demographic Parity–, but in effect it is constrained by a more complex fairness criterion, namely Equalized odds.
First, we proposed an explanation for this difference by looking at the specifics of each algorithm. GridSearch incorporates the fairness constraint as an additional cost inside the cost function of the model. Due to the fairness cost, this family of methods is more likely to deviate from the true labels (increasing false positives), in order to achieve Demographic Parity. On the other hand, FairDream introduces a differentiated attention to harmed individuals inside the cost function of the model. Therefore, it increases accuracy on harmed groups (through enhanced true positives). By basing fairness on accuracy, FairDream-type methods are thus bound to make the overall positive rates converge towards their base rates. As a result, they will rarely achieve Demographic Parity.
Our study of FairDream also led us to a normative conclusion about how to approach fairness. By choosing a conditional metric, the user implicitly commits to the idea that there is sufficient agreement on the transparency of labels and on the shape of their distribution; by selecting an unconditional metric, one commits to the idea that the labels themselves may have introduced unwanted disadvantages, or that the input distribution must be corrected rather than replicated. In the case of FairDream, the default is therefore set to a correction method that may be viewed as overly prudent, but this is as it should be, since as we argued, more revisionary methods need to be carefully justified and cannot be applied lightly. Moreover, we have established a parallel between the correction operated by FairDream and Simpson’s paradox to argue that Equalized Odds expresses a more constrained notion of counterfactual evaluation of fairness than does Demographic Parity.
Our investigation of FairDream allows us to draw a more general lesson concerning the explainability of AI models. Prima facie, FairDream may be seen as operationalizing a simple interventionist fairness metric, namely Demographic Parity. In practice, however, it produces results that are significantly different and that enforce Equalized Odds. As we have argued, the corrections produced by FairDream can be rationalized and justified, but this involves an explanation of the reweighting method used by the algorithm, a contrastive look at nearby systems governed by similar correction mechanisms that produce different outcomes, and a normative justification of the dependence of the correction methods on true labels.
Even an objective as simple in appearance as “equalizing selection percentages across groups” remains largely underdetermined, an aspect that we do not take to be specific to FairDream, but to be shared by other algorithmic intervention methods. We conclude that AI classification systems should preferably be released with an indication of these features: not only which goal they seek to operationalize, but how they actually perform relative to this goal, and then how they perform on alternative data, and relative to minimally modified variants on the same data. Making this information available dispels arbitrariness in classification, and is a step toward more trust, transparency, and control on the side of non-technical users.
Appendix
A FairDream’s Reweighting
We provide here more details on how the sample weights are generated during the in-processing correction of FairDream. The differentiated weights of error are established in a deterministic but simple way, to grant user’s understanding on the process of correction for fairness. For a FairDream model $n$ $\in$ $[1,5]$ of five models to be trained in competition, the new weight of error for each group $S_{k}$} is computed that way:
B Benchmark on Census - models and features
Max Gap between groups OPR FPR TPR ROC-AUC PR-AUC Calibration All
GridSearch 7 9 10 11 12 8 57
FairDream 9 7 6 3 2 8 35
On the same model type and on the same feature, here 16 times, GridSearch and FairDream both attempt to correct the overall selection gap. The 16 data points reported correspond to instances—across model architectures—where a global selection gap of at least 3 : 1 on a variable triggered the corrective phase; FairDream and GridSearch are then compared under these same conditions exposed in Section 5.3. Each “point” in the column of Overall Positive Rates (OPR) in the Table therefore represents a comparison between the best FairDream model and the best GridSearch model on OPR, using the same architecture (e.g. random forest) and the same variable (e.g. sex), on the fairness metric listed at the top of the column. A +1 point of penalty is assigned to FairDream if disparities between the best- and the worst-treated group is at least 5 % above GridSearch; otherwise, no point is given. [^27] This explains why, for metrics such as PR–AUC, the total of the column reaches 14 instead of 16. In Table 3, lower scores indicate better results on that count. To verify the convergence property of FairDream to Equalized Odds, the other columns of the table also indicate the number of times a model was worse than the other at equalizing true positive rates, and false positive rates, across the relevant features.
C Calibration and Precision
@@
Besides Demographic Parity, Equalized Odds is also commonly opposed to various notions of predictive parity, understood either in terms of equal calibration ([Pleiss et al., 2017, Barocas et al., 2023]), or in terms of equal precision relative to a threshold ([Dieterich et al., 2016, Mayson, 2018]). In this appendix, we start with some data relative to FairDream’s performance on calibration, and we explain why, on a normative basis, calibration for us does not constitute a relevant criterion.
@@
One common view of fairness is found in the idea of predictive parity between groups. Predictive parity can mean that the same fraction of true positives is predicted between groups at identical risk scores, or it can mean that the same fraction of true positives is predicted between groups relative to a classification threshold. The first notion corresponds to the statistical notion of calibration of a classifier, whereas the second corresponds to the notion of statistical precision of a classifier. Both notions are sometimes indistinctly referred to under the term “calibration” in the literature (see [Mayson, 2018, Long, 2021]), although strictly speaking they differ. [^28] Both notions have also been invoked in defense of COMPAS against ProPublica’s charges of racial bias. In their response to ProPublica, Northpointe argued that at the cutoff chosen by ProPublica to show Unequal Odds between Whites and Blacks, the statistical precision of the COMPAS algorithm was roughly the same between both groups. Moreover, it was shown that at each risk score, the proportion of rearrest between groups was roughly the same (see [Corbett-Davies et al., 2017, Figure 2]), [Dieterich et al., 2016] produce very closely related data (see their tables A1 and A2), but they look at predictions above all possible risk scores, which is not strictly speaking calibration. this time evidencing calibration between groups.
@@
The upshot is that Equal Precision is neither a necessary, nor a sufficient condition on fairness. In the case we just presented, one might respond that Precision in the “Low” judgments is not constant across the groups, even as precision on the “High” judgments is. But the problem is that to ask for Equal Precision on both labels is tantamount to asking for Equalized Odds, which would undercut Long’s argument. This is not to say that Precision does not matter to FairDream. As explained (see fn. 15), we conducted our benchmark so as to maximize the trade-off between Precision and Recall. From Table 4, we also see that the gap in PR-AUC diminishes in FairDream across groups compared to the baseline. But since this is not a criterion we set a priori, no more than ROC-AUC, we do not build any conclusion on this fact.
@@
In summary, therefore, whether on a descriptive basis, or on a normative basis, Equal Calibration does not end up as a decisive criterion in our evaluation of FairDream’s correction, and neither does Equal Precision.